A man lifts a mass of 3kg upwards through 2m. If g=10m/s2 calculate work done




1
Expert's answer
2021-12-01T10:28:28-0500

"g = 10\frac{m}{s^2}"

"\Delta h = 2m"

"m = 3kg"

"W = mg\Delta h= 3*10*2= 60J"

"\text{Answer :}60\ J"